Sodium metal crystallises in a body centered cubic lattice with a cell age of 425 pm . Calculate the density of the metal.
1
Expert's answer
2019-10-29T06:44:26-0400

Density of unit cell"=\frac{Z\times M}{N_A\times a^3}=0.994\times10^6"